OpenStudy (cheyennecoffelt):

Write the equation of a line that is perpendicular to the given line and that passes through the given point. y-2=7/3 (x+5);(-4,9) You don't have to give me the answer, but steps or a walk through would be nice so i can do it on my own in the future

OpenStudy (michele_laino):

the slope of equation above is: \(m=7/3\) now, the slope \(m'\) of the requested line has to be: \[\Large m' = \frac{{ - 1}}{m} = \frac{{ - 1}}{{\frac{7}{3}}} = ...?\]

OpenStudy (michele_laino):

the requested equation, is: \[y - 9 = m'\left( {x + 4} \right)\]

OpenStudy (cheyennecoffelt):

is it y-9=-3/7 (x-4)?

OpenStudy (cheyennecoffelt):

sorry it took awhile my baby brother came in and started distracting me

OpenStudy (michele_laino):

no problem, here we have to apply this formula: \[y - {y_0} = m'\left( {x - {x_0}} \right)\] where \(x_0=-4\)

OpenStudy (cheyennecoffelt):

y−y0=m′(x−x0) (-4,9) m′=-3/7 applying these gets me y-9=-3/7(x-4) correct? or is it y-9=-3/7(x+4) because you would be subtracting a negative

OpenStudy (cheyennecoffelt):

@Michele_Laino

OpenStudy (michele_laino):

it is the second equation, namely: \[y - 9 = \frac{{ - 3}}{7} \cdot \left( {x + 4} \right)\]

OpenStudy (cheyennecoffelt):

Okay thanks!!! i just wanted to verify so i understand everything.

OpenStudy (michele_laino):

:)